# HiMTok: Learning Hierarchical Mask Tokens for Image Segmentation with Large Multimodal Model

### v1

@@ Abstract
The remarkable performance of large multimodal models (LMMs) has attracted significant interest from the image segmentation community. To align with the next-token-prediction paradigm, current LMM-driven segmentation methods either use object boundary points to represent masks or introduce special segmentation tokens, whose hidden states are decoded by a segmentation model requiring the original image as input. However, these approaches often suffer from inadequate mask representation and complex architectures, limiting the potential of LMMs. In this work, we propose the Hierarchical Mask Tokenizer (HiMTok), which represents segmentation masks with up to 32 tokens and eliminates the need for the original image during mask de-tokenization. HiMTok allows for compact and coarse-to-fine mask representations, aligning well with the LLM next-token-prediction paradigm and facilitating the direct acquisition of segmentation capabilities. We develop a 3-stage training recipe for progressive learning of segmentation and visual capabilities, featuring a hierarchical mask loss for effective coarse-to-fine learning. Additionally, we enable bidirectional information flow, allowing conversion between bounding boxes and mask tokens to fully leverage multi-task training potential. Extensive experiments demonstrate that our method achieves state-of-the-art performance across various segmentation tasks,while also enhancing visual grounding and maintaining overall visual understanding. The codes and models are available at https://github.com/yayafengzi/LMM-HiMTok.
@@ 1 Introduction
There are several paradigms for LLM-based image segmentation. Some works represent a segmentation mask as a sequence of boundary points, which can be learned by a sequence-to-sequence framework [84, 58, 44, 7, 83], as shown in Fig. 2 (a). However, using a limited number of polygon vertices can impede accuracy, particularly in representing masks of complex shapes or multiple regions [44, 84]. Some other LMM-based segmentation methods [27, 49, 76, 47, 78, 2, 64, 61] exploit LLMs to output object hidden states, which are subsequently passed to an additional image-conditioned mask decoder, as shown in Fig. 2 (b). Such hidden states usually correspond to some special learnable tokens (e.g., <SEG> [27]) to adapt to the next-token-prediction paradigm. However, there are three limitations to this paradigm. First, due to the reliance on powerful image segmentation models, large language models (LLMs) have insufficient learning of precise spatial localization in images. Second, there is an inconsistency in mask representation between LLM input and output. Special tokens serve solely as identifiers and are used as LLM input, resulting in the loss of crucial information from the corresponding hidden states in autoregressive modeling. If users provide mask prompts, they need to extract input prompt features through RoI pooling or cross-attention with visual features [78, 76]. Third, the overall architecture is complex and heavy. The mask decoder is typically designed based on previous segmentation models and requires the original image to be used again, such as SAM [26] and Mask2Former [14], with some even requiring an additional vision encoder. Overall, these limitations restrict LMMs from achieving their full potential in image segmentation.
@@
In this work, we propose HiMTok, an effective Hierarchical Mask Tokenizer that can represent a mask image using up to 32 hierarchical tokens. Inspired by TiTok [74], images can be expressed as highly compact 1D sequences. Based on this concept, we introduce a hierarchical mask tokenizer that uses hierarchical mask loss and a causal attention mechanism to represent mask images from coarse to fine. In this representation, earlier tokens mainly correspond to coarse locations and prototypes, while later tokens focus more on local fine-grained details. Each mask token is tightly conditioned on the preceding ones. Hence, this hierarchical design aligns seamlessly with the autoregressive principle of LLMs. By considering mask tokens as a new language, it enables LMMs to conveniently gain image segmentation capabilities. As shown in Fig. 2 (c), LMM generates mask tokens just as the way for text tokens, and the final segmentation mask is obtained by de-tokenizing the mask tokens, without the need for the input image.
@@
Through extensive experiments, our method achieves state-of-the-art performance on image segmentation tasks, including referring expression segmentation, reasoning segmentation and open-vocabulary segmentation. Our model also improves referring expression comprehension, a visual grounding task. These results demonstrate the effectiveness of HiMTok for multi-task learning in LMM. Meanwhile, our competitive performance in general image understanding tasks indicates that HiMTok enables LMMs to acquire segmentation capabilities without compromising the general abilities.
@@
We propose HiMTok, an efficient hierarchical mask tokenizer capable of representing a mask image using up to 32 hierarchical tokens. LMMs are able to learn such sequences conveniently for effective image segmentation, without the need for an image-conditioned mask decoder or off-the-shelf segmentation foundation models.
@@
Extensive experiments demonstrate that the LMM equipped with HiMTok not only shows superiority in various image segmentation tasks, but also improves visual grounding and maintains the general image understanding capability.
@@ 2 Related Works
Large multimodal models (LMMs) have recently become a popular topic. We have witnessed the continuous improvement in tasks such as visual understanding, OCR, and visual grounding [62, 11, 69, 57, 16, 53, 30]. However, as another fundamental visual task, image segmentation has not been widely included in the capabilities of these open-source large models. In this section, we will review works on image segmentation paradigms in Fig. 2 (a) (b) and image tokenization, to better understand our method.
@@ 3.1 Overview
To this end, we design a three-stage training recipe for progressive and efficient learning (Fig. 3). HiMTok is first fully trained to enable the single-modality mask tokenization and de-tokenization. In stage 2, the joint training of LMM and HiMTok is performed, which brings vision-language alignment in mask tokenization, while still being efficient with images of low resolution. In stage 3, we focus solely on finetuning the LMM using high-resolution images to optimize performance.
@@ 3.2 Hierarchical Mask Tokenizer
HiMTok consists of three components: a mask tokenizer (MT), a vector quantization layer (VQ), and a mask de-tokenizer (MD), as illustrated in Fig. 3 (a). Following TiTok [74], some 1D learnable latent tokens are encoded by MT with mask image patches. After quantized by a VQ variant [85], the encoded latent tokens are then reconstructed by MD. The number of 1D mask tokens is much smaller than the number of 2D mask patches.
@@
Hierarchical mask loss. Explicit supervision on mask tokens of different levels is necessary to ensure hierarchy. At the -th hierarchical level, the first mask tokens are de-tokenized to independently. In the full level, all mask tokens are used. Based on our observation, few tokens (e.g., 4) usually lead to coarse Gaussian distributions appearing in the mask image after the mask de-tokenizer. Thus, we empirically exploit the Gaussian blur with different sizes of Gaussian kernel on different levels for multi-grained mask label () preparation. Details on how we make coarse mask labels are illustrated in Appendix A. The hierarchical mask loss (HML) is calculated across the different levels:
@@ 3.3 Mask-aware Large Multimodal Model
By regarding the mask tokens in HiMTok as a new language, we can easily incorporate image segmentation capabilities into LMMs. The only modification required for LMM is an expanded vocabulary with the mask token codebook.
Prompt format. To highlight and distinguish the segmentation task and mask tokens from natural language, we exploit some special tokens to indicate them. Simply put, we use “<ref>referred object</ref>” to pinpoint the object to be segmented, and <mt_start>mt_i <mt_end> to indicate the segmentation mask tokens. Besides, visual grounding is also involved in our experiments as a regular vision-language task, where we use “<box>[x1,y1,x2,y2]</box>” to indicate the bounding box. More details about the prompt design are listed in Appendix G.
@@
Progressive training. To enable LMM to learn mask tokens while maintaining its general capability, we train LMM in two stages, as shown in Fig. 3 (b) (c). In stage 2, a massive amount of image segmentation data is used along with some general data. The mask tokenization is aligned with vision-language through joint training. Since the LMM encounters various segmentation cases, the understanding of mask tokens is gradually improved. Low-resolution images are used to train effectively. Both the cross entropy loss for next-token-prediction and the hierarchical mask loss are adopted for optimization. In stage 3, the LMM is further trained with high-resolution images as input. The amount of segmentation data is reduced to avoid collapse on the general capacity. Only the cross entropy loss is used.
Inference. During inference, if mask tokens are detected in the LLM output, the lightweight mask de-tokenizer will visualize them as the final predicted segmentation mask. Our method achieves its performance independently of segmentation foundation models, yet they can still be considered for further potential enhancement. Flexibly, the mask from our de-tokenizer can be further refined by feeding it into an independently finetuned SAM [26]. Details on SAM finetuning for optional usage can be found in Appendix B. We do not utilize SAM in our experiments unless explicitly stated.
@@ 4.1 Training Datasets
Since the LMM is expected to learn brand-new mask tokens for image segmentation while maintaining its original capability, both segmentation data and general-purpose data are required for effective training. All datasets transcribed in the 3 stages are listed in Tab. 1. For stage-1 training, we exploit the groundtruth masks in all the segmentation-related datasets to train our HiMTok. Random cropping is adopted to augment the binary mask images. For stage 2, all the listed datasets are used with a total amount of 7.1 million. The ratio of segmentation data reaches up to 0.41, which accelerates the learning of mask tokens. For stage 3, we discard a large number of segmentation and detection samples that are low-quality or overly simplistic in instruction. The details of the down-sampled datasets are listed in Appendix C. As a result, the total amount of training data is decreased to 5.0 million, with the ratio of segmentation data dropping to 0.24, which ensure the general performance.
@@ 4.4 Reasoning Segmentation
As shown in Tab. 4, our method achieves the best performance with the previous SOTA of similar model size in ReasonSeg [27]. Notably, our scores on both the validation and test sets are nearly identical, whereas other methods show lower performance on the test set, which contains a large proportion of short and sophisticated questions. This further demonstrates that our model not only possesses strong segmentation capabilities but also retains powerful text understanding.
@@ 4.6 Referring Expression Comprehension
Here we present the results on the REC benchmarks, which are classic in visual grounding. Our output boxes are generated directly by the LMM following the mask tokens, without any post-processing based on segmentation masks. This approach is highly efficient if our goal is solely object detection, as it bypasses the de-tokenization process.
@@
4.7 General Image Understanding
We compare our model with the baseline on MME [19]. As shown in Tab. 7, our model is comparable to InternVL2.5-8B across various dimensions. The joint learning of segmentation even brings some promotion in image understanding of some areas, e.g., position, scene. However, the performance drops in some dimensions, which may be due to the lack of diverse and high-quality data in this segmentation-oriented version. For example, we do not incorporate data on celebrity and OCR. More results on the general image understanding are listed in Appendix E.
@@ 4.8.1 Effect of the mask token length
Considering that our model was trained with full-length ( in Eq. 1) mask token sequences, we finetune it using a mixture of token sequences with varying lengths, by specifying the length in the prompt. Details on the prompt format can be found in Appendix G. We evaluate on the RefCOCO validation set.
@@
For REC, the object bounding box is predicted after mask tokens of varying lengths. We report accuracy on 3 IoU thresholds: 0.5, 0.7 and 0.9, which correspond to evaluations of varying levels from coarse to fine. Fig. 5 shows the effect on REC. As the role segmentation plays, more mask tokens contribute to more refined box prediction. As more mask tokens are used, acc@0.5 sees a slight improvement, while acc@0.9 improves significantly. These findings prove the information flow from mask tokens to box coordinates is very useful in both training and evaluation phase. From some perspective, we can regard the hierarchical mask tokens as a chain-of-thought for visual grounding. By the way, we do not see improvement on RES for the information flow from box to mask. This may be due to the fact that the former mask tokens are easier to generate than the direct box coordinates.
@@ 4.8.2 Effect of hierarchical mask loss
As shown on the left of Fig. 6, mask tokens fewer than 32 lead to completely incorrect segmentation masks, which suggests that training without HML strictly requires the use of full-length mask tokens throughout. In contrast, the use of our hierarchical mask tokens is flexible. More interesting visualization cases are presented in Appendix F.
@@ 5 Conclusion
Appendix A Multi-grained mask labels
@@ Appendix B Finetuning SAM
Appendix C More details on training
In stage-3, some datasets are down-sampled. Details are listed in Tab. 9. The initial learning rate is in stage 2, in stage 3, and in task finetuning.
@@ Appendix C More details on training
Appendix D Mask Perception
One key feature of our method is the consistence of mask representation in LLM input and output. The image segmentation tasks are all text-to-mask. As a supplement, we devise a mask-to-text task, Mask Perception, which aims for fine-grained understanding. Mask Perception (MaP) mirrors RES: models are required to choose a matched expression, given an image, an object mask and several expression options.
The training and test sets are built based on RefCOCO/+/g [72, 40]***The Mask Perception dataset will be released publicly. We randomly sample some examples and reverse the positions of masks and expressions in multiple-choice questions. One or more positive options are possible in the training set, while only one positive option is available in the test set to simplify evaluation. Negative expression options are selected from other objects in the same image or from different images. This approach encourages the model to distinguish the distinctive features of various parts. Statistics are shown in Tab. 10.

### v2

@@ Abstract
The remarkable performance of large multimodal models (LMMs) has attracted significant interest from the image segmentation community. To align with the next-token-prediction paradigm, current LMM-driven segmentation methods either use object boundary points to represent masks or introduce special segmentation tokens, whose hidden states are decoded by a segmentation model requiring the original image as input. However, these approaches often suffer from inadequate mask representation and complex architectures, limiting the potential of LMMs. In this work, we propose the Hierarchical Mask Tokenizer (HiMTok), which represents segmentation masks with up to 32 tokens and eliminates the need for the original image during mask de-tokenization. HiMTok allows for compact and coarse-to-fine mask representations, aligning well with the LLM next-token-prediction paradigm and facilitating the direct acquisition of segmentation capabilities. We develop a 3-stage training recipe for progressive learning of segmentation and visual capabilities, featuring a hierarchical mask loss for effective coarse-to-fine learning. Additionally, we enable bidirectional information flow, allowing conversion between bounding boxes and mask tokens to fully leverage multi-task training potential. Extensive experiments demonstrate that our method achieves state-of-the-art performance across various segmentation tasks,while also enhancing visual grounding and maintaining overall visual understanding. The codes are available at https://github.com/yayafengzi/LMM-HiMTok.
@@ 1 Introduction
There are several paradigms for LLM-based image segmentation. Some works represent a segmentation mask as a sequence of boundary points, which can be learned by a sequence-to-sequence framework [84, 58, 44, 7, 83], as shown in Fig. 2 (a). However, using a limited number of polygon vertices can impede accuracy, particularly in representing masks of complex shapes or multiple regions [44, 84]. Some other LMM-based segmentation methods [27, 49, 76, 47, 78, 2, 64, 61] exploit LLMs to output object hidden states, which are subsequently passed to an additional image-conditioned mask decoder, as shown in Fig. 2 (b). Such hidden states usually correspond to some special learnable tokens (e.g., <SEG> [27]) to adapt to the next-token-prediction paradigm. However, there are three limitations to this paradigm. First, due to the reliance on powerful image segmentation models, large language models (LLMs) have insufficient learning of precise spatial localization in images. Second, there is an inconsistency in mask representation between LLM input and output. Special tokens serve solely as identifiers and are used as LLM input, resulting in the loss of crucial information from the corresponding hidden states in autoregressive modeling. If users provide mask prompts, they need to extract input prompt features through RoI pooling or cross-attention with visual features [78, 76]. Third, the overall architecture is complex and heavy. The mask decoder is typically designed on previous segmentation models and requires the original image to be used again, such as SAM [26] and Mask2Former [14], with some even requiring an additional vision encoder. Overall, these limitations restrict LMMs from achieving their full potential in image segmentation.
@@
In this work, we propose HiMTok, an effective Hierarchical Mask Tokenizer that can represent a mask image using up to 32 hierarchical tokens. Inspired by TiTok [74], images can be expressed as highly compact 1D sequences. Based on this concept, we introduce a hierarchical mask tokenizer that uses hierarchical mask loss and a causal attention mechanism to represent mask images from coarse to fine, which can be learned from abundant and easily accessible mask data. In this representation, earlier tokens mainly correspond to coarse locations and prototypes, while later tokens focus more on local fine-grained details. Each mask token is tightly conditioned on the preceding ones. Hence, this hierarchical design aligns seamlessly with the autoregressive principle of LLMs. By considering mask tokens as a new language, it enables LMMs to conveniently gain native segmentation capabilities without external segmentation foundation models. As shown in Fig. 2 (c), LMM generates discrete mask tokens just as the way for text tokens, and the final segmentation mask is obtained by de-tokenizing the mask tokens.
@@
Through extensive experiments, our method achieves state-of-the-art performance on image segmentation tasks, including referring expression segmentation, reasoning segmentation and open-vocabulary segmentation. Our model also improves referring expression comprehension, a visual grounding task. These results demonstrate the effectiveness of our elegant HiMTok for multi-task learning in LMM. Meanwhile, our competitive performance in general image understanding tasks indicates that HiMTok enables LMMs to acquire segmentation capabilities without compromising the general abilities.
@@
We propose HiMTok, an efficient hierarchical mask tokenizer capable of representing a mask image using up to 32 hierarchical tokens. LMMs are able to learn such token sequences for effective image segmentation, without the need for an image-conditioned mask decoder or off-the-shelf segmentation foundation models.
@@
Extensive experiments demonstrate that LMM equipped with HiMTok not only shows superiority in various image segmentation tasks, but also improves visual grounding and maintains the general image understanding capability. Interestingly, visual chain-of-thought by “outputting mask tokens before box” improves visual grounding.
@@ 2 Related Works
Large multimodal models (LMMs) have recently become a popular topic. We have witnessed the continuous improvement in tasks such as visual understanding, OCR, and visual grounding [62, 11, 69, 57, 16, 53, 30]. However, as another fundamental visual task, image segmentation [86] has not been widely included in the capabilities of these open-source large models. In this section, we will review works on image segmentation paradigms in Fig. 2 (a) (b) and image tokenization, to better understand our method.
@@ 3.1 Overview
To this end, we design a three-stage training recipe for progressive and efficient learning (Fig. 3). HiMTok is first fully trained to enable the single-modality mask tokenization and de-tokenization. In stage 2, the joint training of LMM and HiMTok is performed, which brings vision-language alignment in mask tokenization, while still being efficient with images of low resolution. In stage 3, we focus solely on finetuning the LMM with high-resolution images.
@@ 3.2 Hierarchical Mask Tokenizer
HiMTok consists of three components: a mask tokenizer (MT), a vector quantization layer (VQ), and a mask de-tokenizer (MD), as illustrated in Fig. 3 (a). Following TiTok [74], some 1D learnable latent tokens are encoded by MT along with mask image patches. After quantized by a VQ variant [85], the encoded latent tokens are then reconstructed by MD***MT annd MD are both standard Transformer layers. The number of 1D mask tokens is much smaller than the number of 2D mask patches.
@@
Hierarchical mask loss. Explicit supervision on mask tokens of different levels is necessary to ensure hierarchy. At the -th hierarchical level, the first mask tokens are de-tokenized to independently. In the full level, all mask tokens are used. Based on our observation, few tokens (e.g., 4) usually lead to coarse Gaussian distributions appearing in the mask image after the mask de-tokenizer. Thus, we empirically exploit the Gaussian blur with different sizes of Gaussian kernel on different levels for multi-grained mask label () preparation. Details on how we make coarse mask labels are illustrated in Appendix B. The hierarchical mask loss (HML) is calculated across the different levels:
@@ 3.3 Mask-aware Large Multimodal Model
Regarding the mask tokens in HiMTok as a new language, we can easily incorporate image segmentation capabilities into LMMs. The only modification required for LMM is an expanded vocabulary with the mask token codebook.
Prompt format. To highlight and distinguish the segmentation task and mask tokens from natural language, we exploit some special tokens to indicate them. Simply put, we use “<ref>referred object</ref>” to pinpoint the object to be segmented, and <mt_start>mt_i <mt_end> to indicate the segmentation mask tokens. Besides, visual grounding is also involved in our experiments as a regular vision-language task, where we use “<box>[x1,y1,x2,y2]</box>” to indicate the bounding box. More details about the prompt design are listed in Appendix I.
@@
Progressive training. To enable LMM to learn mask tokens while maintaining its general capability, we train LMM in two stages, as shown in Fig. 3 (b) (c). In stage 2, a massive amount of image segmentation data is used along with some general data. The mask tokenization is aligned with vision-language through joint training. Since the LMM encounters various segmentation cases, the understanding of mask tokens is gradually improved. Low-resolution images are used to train effectively. Both the cross entropy loss () for next-token-prediction and the hierarchical mask loss are adopted for optimization. In stage 3, the LMM is further trained with high-resolution images as input. The amount of segmentation data is reduced to avoid collapse on the general capacity. Only the cross entropy loss is used.
Inference. During inference, if mask tokens are detected in the LLM output, the lightweight mask de-tokenizer will visualize them as the final predicted segmentation mask. Alternatively, the mask from our de-tokenizer can be further potentially refined by feeding it into an independently finetuned SAM [26]. Details on SAM finetuning for optional usage can be found in Appendix C. We do not utilize SAM in our experiments by default, unless explicitly stated.
@@ 4.1 Training Datasets
Since the LMM is expected to learn brand-new mask tokens for image segmentation while maintaining its original capability, both segmentation data and general-purpose data are necessary for effective training. All datasets transcribed in the 3 stages are listed in Tab. 1. For stage-1 training, we exploit the groundtruth masks in all the segmentation-related datasets to train our HiMTok. Random cropping is adopted to augment the binary mask images. For stage 2, all the listed datasets are used with a total amount of 7.1 million. The ratio of segmentation data reaches up to 0.41, which accelerates the learning of mask tokens. For stage 3, we discard a large number of segmentation and detection samples that are low-quality or overly simplistic in instruction. The details of the down-sampled datasets are listed in Appendix D. As a result, the total amount of training data is decreased to 5.0 million, with the ratio of segmentation data dropping to 0.24, which ensure the general performance.
@@ 4.4 Reasoning Segmentation
As shown in Tab. 4, our method achieves the best performance with the previous SOTA of similar model size in ReasonSeg [27], a benchmark featuring complicated and obscure instructions. Notably, our scores on both the validation and test sets are nearly identical, whereas other methods show lower performance on the test set, which contains a large proportion of short and sophisticated questions. This further demonstrates that our model not only possesses strong segmentation capabilities but also retains powerful text understanding.
@@ 4.6 Referring Expression Comprehension
Here we present the results on the REC benchmarks, which are widely recognized for visual grounding evaluation. Our output boxes are generated directly by the LMM following the mask tokens, without any post-processing based on segmentation masks. This approach is highly efficient if our goal is solely object detection, as it bypasses the de-tokenization process.
@@
4.7 General Image Understanding
We compare our model with the baseline on MME [19]. As in Tab. 7, our model is comparable to InternVL2.5-8B across various dimensions. The joint learning of segmentation even brings some promotion in image understanding of some areas, e.g., position, scene. However, the performance drops in some dimensions, which is due to the lack of diverse and high-quality data in this segmentation-oriented version. For example, we do not specially incorporate data on celebrity and OCR. This also explains why the performance of LMM(ft) differs significantly across the dimensions: the segmentation data we used has stronger bias on existence and scene than others. The model retains intrinsic semantic understanding, despite being fine-tuned exclusively on segmentation data. More results on the general image understanding are listed in Appendix G.
@@ 4.8.1 Effect of the mask token length
Considering that our model was trained with full-length ( in Eq. 1) mask token sequences, we finetune it using a mixture of token sequences with varying lengths, by specifying the length in the prompt. Details on the prompt format can be found in Appendix I. We evaluate on the RefCOCO validation set.
@@
For REC, the object bounding box is predicted after mask tokens of varying lengths. It is exactly a kind of visual chain-of-thought. We report accuracy on 3 IoU thresholds: 0.5, 0.7 and 0.9, which correspond to evaluations of varying levels from coarse to fine. Fig. 5 shows the effect on REC. As the role segmentation plays, more mask tokens contribute to more refined box prediction. As more mask tokens are used, acc@0.5 sees a slight improvement, while acc@0.9 improves significantly. These findings prove the information flow from mask tokens to box coordinates is very useful in both training and evaluation phase. From some perspective, we can regard the hierarchical mask tokens as a chain-of-thought for visual grounding. By the way, we do not see improvement on RES for the information flow from box to mask. This may be due to the fact that the former mask tokens are easier to generate than the direct box coordinates.
@@ 4.8.2 Effect of hierarchical mask loss
As shown on the left of Fig. 6, mask tokens fewer than 32 lead to completely incorrect segmentation masks, which suggests that training without HML strictly requires the use of full-length mask tokens throughout. In contrast, the use of our hierarchical mask tokens is flexible. More interesting visualization cases are presented in Appendix H.
@@ 5 Conclusion
Appendix A Limitations
While HiMTok enables large multimodal models to acquire native (referring) image segmentation capabilities in a concise and natural manner, the current work has some limitations. (1) The length of predicted mask tokens is pre-defined. LMMs are not able to determine it adaptively according to object shape complexities. (2) The current model is relatively passive for object segmentation, due to the use of passive segmentation training data. We need to specify referring expressions to clarify the expected objects, rather than let the model itself segment all objects of interest at once. (3) It appears challenging for fine-grained region segmentation in the current version (See Appendix F). The lack of multi-scale feature design may cause the loss of fine-grained features.
Appendix B Multi-grained mask labels
@@ Appendix C Finetuning SAM
Appendix D More details on training
In stage-3, some datasets are down-sampled. Details are listed in Tab. 9. The initial learning rate is in stage 2, in stage 3, and in task finetuning. The GPU hours (Nvidia A800) for our 3 stages are: 192, 1920 and 640.
@@ Appendix D More details on training
Appendix E Mask Perception
One key feature of our method is the consistence of mask representation in LLM input and output. The image segmentation tasks are all text-to-mask. As a supplement, we devise a mask-to-text task, Mask Perception†††https://huggingface.co/datasets/yayafengzi/Mask_Perception, which aims for fine-grained understanding. Mask Perception (MaP) mirrors RES: models are required to choose a matched expression, given an image, an object mask and several expression options.
The training and test sets are built based on RefCOCO/+/g [72, 40]. We randomly sample some examples and reverse the positions of masks and expressions in multiple-choice questions. One or more positive options are possible in the training set, while only one positive option is available in the test set to simplify evaluation. Negative expression options are selected from other objects in the same image or from different images. This approach encourages the model to distinguish the distinctive features of various parts. Statistics are shown in Tab. 10.